# Tactile-Morph Skills: Energy-Based Control Meets Data-Driven Learning

### v1

@@ 3 Methodology
The Tactile-Morph Skills framework is designed to achieve three primary objectives: enabling a system to i.) “understand” its environment, adapt the prior policy, and learn the correct control for varying environmental conditions, and ii.) robustly and safely adjust to unforeseen process conditions, as illustrated in Fig. 1. To achieve these goals, we integrated a Temporal Convolutional Network (TCN) to estimate the energy distribution of manipulation tasks, allowing precise management of the energy tank status. In the following sections, we introduce the details of the nominal skill, the TCN-based energy status estimation, and the overall structure of our proposed Tactile-Morph skills framework.
Figure 1: Pipeline Overview: Tactile-Morph Skills Framework is composed of a energy distribution estimator and unified force-impedance control(UFIC). During the training phase, a set of tactile skills, designed based on basic motion patterns, is planned using point cloud observations of the working surface. These tactile skills are executed by the robot, and the corresponding energy distribution is recorded in real-time. The collected data is then utilized to train our energy distribution estimation model. Upon completion of the training process, the energy distribution for any novel tactile skill sampled from the working surface can be accurately estimated. By integrating the estimated energy distribution with UFIC, we can provide a formal guarantee of passivity to any contact-rich tactile skill.
@@ 3.3 Pipeline Overview
Using the trained model, we estimate the energy needed for any randomly generated tactile skill starting at a specific position on the working surface. An energy heat map is constructed by leveraging the outputs of energy estimation and linear interpolation, as shown in Fig. 3. This energy heat map enables fast queries of the required energy trajectory for any tactile skill on the working surface, allowing for robust recovery, extension, and re-planning in real time.
@@ 4.1 Performance Evaluation
Methods MSE MAPE(%) $\text{MAPE}_{\text{sum}}$(%)

### v2

@@ 3 Methodology
The Tactile-Morph Skills framework is designed to achieve three primary objectives: enabling a system to i.) ”understand” its environment, adapt the prior policy, and learn the correct control for varying environmental conditions, and ii.) robustly and safely adjust to unforeseen process conditions, as illustrated in Fig. 1. To achieve these goals, we integrated a Temporal Convolutional Network (TCN) to estimate the energy distribution of manipulation tasks, allowing precise management of the energy tank status. In the following sections, we introduce the details of the nominal skill, the TCN-based energy status estimation, and the overall structure of our proposed Tactile-Morph skills framework.
Figure 1: Pipeline Overview: Tactile-Morph Skills Framework is composed of an energy distribution estimator and unified force-impedance control(UFIC). During the training phase, a set of tactile skills, designed based on basic motion patterns, is planned using point cloud observations of the working surface. These tactile skills are executed by the robot, and the corresponding energy distribution is recorded in real-time. The collected data is then utilized to train our energy distribution estimation model. Upon completion of the training process, the energy distribution for any novel tactile skill sampled from the working surface can be accurately estimated. By integrating the estimated energy distribution with UFIC, we can provide a formal guarantee of passivity to any contact-rich tactile skill.
@@ 3.3 Pipeline Overview
Using the trained model, we estimate the energy needed for any randomly generated tactile skill starting at a specific position on the working surface. An energy heat map is constructed by leveraging the outputs of energy estimation and linear interpolation, as shown in Fig. 3. This energy heat map enables fast queries of the required energy trajectory for any tactile skill on the working surface, allowing for robust recovery, extension, and re-planning in real-time.
@@ 4.1 Performance Evaluation
Methods MSE MAPE(%) $\mathrm{MAPE}_{sum}(\%)$